• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Mina Kat

    Log entry time 00:03:22 on April 8,2007

    Entry number 198134

    keyword=Shift Summary

    16:20 MCC Called will try to give us 80uA.
    Asked to stay at 70 uA see if everything is ok with target and
    then go to max current.
    16:30 HPH too low, adjusted JT valve. Flow now ~20 g/s

    16:48 Asked MCC to ramp up current from 70 to 80 uA in steps of 2 uA
    while checking target cooling system.

    Right spectrometer Q1 trips, power supply window indicates "Earth Leakage Failure".
    17:15 Paged Ed Folts, requested Hall access to cycle power supplies.
    Q1 trips two more times, requested second Hall access.
    Ed Folts, Jack Segal looking into finding and fixing problem.

    19:30 Started monitoring behavior of Q1 (current, flow etc) with strip tool while a gentleman from Rad Con is surveying the Hall.
    19:37 Hall survey completed.
    20:26 Ed Folts back in Hall to increase all quad flows.
    20:35 Ed back from Hall, requested beam
    20:56 Beam back at 70uA
    Downloaded trigger, got no error message
    Run 2167: short test run to check beam spot and prescale factors
    Start production runs at Q2= 25 fm-2:
    Runs 2168-2174, at current 71 uA
    Run 2175 dummy target at 71 uA